OpenStudy (anonymous):

A line is drawn through points (0,6) and (3,0) write an equation that fits the line

OpenStudy (neer2890):

equation of a line passing from two given points (x1,y1) and (x2,y2) is \[y-y1=(y2-y1)\frac{ x-x1 }{ x2-x1 }\]

OpenStudy (neer2890):

Put all the values of x1,x2,y1 and y2 you will get the equation of a line as 2x+y=6